# Constructing Per-Shot Bitrate Ladders using Visual Information Fidelity

### v1

@@ V-D Predicting Quality or Bitrate using VIF Features
V-E Predicting Quality or Bitrate using an Ensemble of Low-Level features and VIF features
@@
Fig. 6 shows the performances of the resulting learned bitrate and quality prediction models. It may be observed that the quality prediction models were slightly more successful (better correlations) than the bitrate prediction models. The regressors trained on $\text{LLF}_{2}$/$\text{LLF}_{3}$ and $\text{VIFF}_{3}/\text{VIFF}_{6}/\text{VIFF}_{9}$ yield slightly better performance than the other models across resolutions. The performances of these models was similar to the best performing models trained on either low-level features or VIF features. However, it should be understood that these correlations do not indicate the potential performance gains of these models when used to construct bitrate or quality ladders, or their BD-Rate or BD-VMAF gains, against fixed bitrate ladders.
@@
V-F Correction Algorithms
In the preceding, we have studied various techniques for predicting per-shot bitrate and quality ladders, using video quality and bitrate prediction models, respectively. Given that machine learning models are prone to errors, it is important to consider corrective measures on the predicted ladders to rectify any inaccuracies. In [8], we suggested a correction mechanism such that traversing the bitrate ladder from top to bottom (from higher to lower bitrates), will satisfy the condition that the optimal resolution at a current bitrate step should no larger than the optimal resolution at the previous step. This strategy works because:
@@ V-F Correction Algorithms
  • When constructing a bitrate ladder, we use the predicted quality scores of compressed videos at each bitrate, but over all resolutions, whereby the resolution having the highest predicted quality is selected as the optimal resolution for encoding at that bitrate.
@@
Tables VII and VIII depict simulated examples of these Top-to-Bottom and Bottom-to-Top ladder correction strategies applied to bitrate ladders and quality ladders, respectively. These correction techniques are applied after the bitrate or quality ladders are constructed.
@@
(b) After Correction
TABLE VII: An example of Top-to-Bottom bitrate ladder correction.
@@
(b) After Correction
TABLE VIII: An example of Bottom-to-Top quality ladder correction.
V-G Bjontegaard Delta Metrics
One of the most commonly used techniques for evaluating or comparing the compression efficiency of video codecs, presets, or encoding modes is by calculating bitrate and quality savings using the Bjøntegaard Delta (BD) method [30]. It measures the average difference between two RQ curves by fitting a third-order cubic polynomial through the data points, and then calculating the integral to estimate the average bitrate savings, referred to as BD-Rate, and the average quality savings, referred to as BD-Quality (e.g., BD-PSNR or BD-VMAF). Although there are many implementations of this concept, based on the results in [31], we used the implementation proposed in [32].
Since we are performing constant-quality encoding, at each step $b_{i}$ in the bitrate ladder and its corresponding optimal resolution $R_{i}$, to construct the convex hull, we utilize the RQ points of the video from the dataset having resolution of $R_{i}$, with bitrates lying within the range [$b_{i}$, $b_{i+1}$]. We use this process, rather than selecting the nearest point, to obtain a precise convex hull for BD metric calculations. We applied a similar procedure when constructing convex hulls for quality ladders. We calculated the BD-Rate and BD-VMAF on each video in the validation and test datasets against the fixed bitrate ladder [3] and the reference bitrate ladder. The reference bitrate ladders were constructed by sampling the convex hull constructed by exhaustive encoding using the bitrate ladder steps we previously mentioned. We report the mean and standard deviation of BD-Rate and BD-VMAF for each of the compared techniques, on the combined validation and test datasets.
@@ V-G Bjontegaard Delta Metrics
V-H Constructing Per-Shot Bitrate and Quality Ladders
The construction of per-shot reference bitrate ladders is achieved by sampling the convex hulls computed on videos from both the validation and test datasets. The reference bitrate ladder demonstrated a BD-Rate performance having a mean improvement of $-20.63\%$ (negative BD-rate means a gain in bitrate) and a standard deviation of $17.124\%$, and a BD-VMAF performance having a mean improvement of $4.473$ (positive means gain in quality) and standard deviation $4.139$ against the fixed bitrate ladder [3]. Our models were compared against techniques that predict cross-over bitrates or cross-over VMAF points. For more accurate comparisons, we computed the performance of bitrate and quality ladders constructed using true cross-over bitrates and VMAFs points, respectively. We found that the bitrate ladder constructed using true cross-over bitrates, attained a BD-rate performance having a mean improvement of $-21.653\%$, and a standard deviation of $16.835\%$. It also obtained a BD-VMAF performance having a mean improvement of $4.712$ and a standard deviation of $4.091$. In a similar vein, the quality ladder constructed using true cross-over VMAF points, demonstrated a BD-rate performance having a mean improvement of $-21.847\%$ and a standard deviation of $16.254\%$, and a BD-VMAF performance having a mean improvement of $4.752$ and standard deviation $4.145$.
TABLE IX: Means and standard deviations of BD-metrics, and closeness of each model’s predicted per-shot bitrate ladders against fixed and reference bitrate ladders on the validation and test datasets.
@@ V-H Constructing Per-Shot Bitrate and Quality Ladders
$\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-9.256/20.964$ $2.084/4.369$ $14.053/11.451$ $-2.144/1.933$ $0.5$ $0.433$ $0.167$
$\text{LLF}_{2}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.429/16.747$ $4.323/4.092$ $\mathbf{0.439}/4.791$ $\mathbf{-0.016}/1.004$ $0.967$ $0.933$ $0.767$
$\text{VIFF}_{1}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-12.238/22.332$ $2.79/4.704$ $9.78/13.017$ $-1.519/2.145$ $0.7$ $0.6$ $0.4$
$\text{VIFF}_{2}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-16.022/20.536$ $3.521/4.678$ $5.381/8.955$ $-0.922/1.617$ $0.767$ $0.767$ $0.633$
$\text{VIFF}_{3}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-17.372/18.17$ $3.665/4.184$ $3.746/5.768$ $-0.628/1.039$ $0.767$ $0.7$ $0.567$
$\text{VIFF}_{4}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-14.537/19.764$ $3.258/4.506$ $7.153/8.783$ $-1.141/1.584$ $0.733$ $0.6$ $0.533$
$\text{VIFF}_{5}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-16.255/19.114$ $3.476/4.282$ $4.907/7.464$ $-0.841/1.319$ $0.767$ $0.667$ $0.533$
$\text{VIFF}_{6}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-17.779/18.186$ $3.758/4.218$ $3.757/5.967$ $-0.629/1.065$ $0.833$ $0.7$ $0.533$
$\text{VIFF}_{7}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-14.591/20.166$ $3.252/4.508$ $7.147/9.862$ $-1.087/1.554$ $0.7$ $0.667$ $0.5$
$\text{VIFF}_{8}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-16.362/18.327$ $3.416/4.134$ $4.954/7.01$ $-0.806/1.191$ $0.8$ $0.7$ $0.5$
$\text{VIFF}_{9}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-17.889/17.545$ $3.828/4.11$ $3.103/5.172$ $-0.519/0.904$ $0.867$ $0.733$ $0.567$
$\text{LLF}_{2}$, $\text{VIFF}_{1}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.151/16.703$ $4.195/3.889$ $1.025/4.703$ $-0.086/0.933$ $0.967$ $0.867$ $0.733$
$\text{LLF}_{2}$, $\text{VIFF}_{2}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.523/17.207$ $4.39/4.071$ $0.829/5.388$ $-0.065/1.073$ $0.967$ $0.867$ $0.767$
$\text{LLF}_{2}$, $\text{VIFF}_{3}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.201/16.083$ $4.259/3.85$ $1.113/5.015$ $-0.089/0.943$ $1.0$ $0.9$ $0.8$
$\text{LLF}_{2}$, $\text{VIFF}_{4}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.061/16.597$ $4.361/4.089$ $1.587/6.181$ $-0.142/1.045$ $0.933$ $0.8$ $0.7$
$\text{LLF}_{2}$, $\text{VIFF}_{5}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-19.597/16.585$ $4.115/3.909$ $2.106/5.344$ $-0.268/0.891$ $0.933$ $0.867$ $0.667$
$\text{LLF}_{2}$, $\text{VIFF}_{6}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.495/16.23$ $4.405/3.878$ $1.073/4.856$ $-0.072/0.936$ $1.0$ $0.933$ $0.833$
$\text{LLF}_{2}$, $\text{VIFF}_{7}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $\mathbf{-20.761}/16.528$ $\mathbf{4.41}/3.898$ $0.592/4.172$ $-0.017/0.803$ $0.933$ $0.9$ $0.867$
$\text{LLF}_{2}$, $\text{VIFF}_{8}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.516/15.94$ $4.296/3.838$ $0.767/5.013$ $-0.018/0.89$ $1.0$ $0.9$ $0.8$
$\text{LLF}_{2}$, $\text{VIFF}_{9}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-19.953/16.822$ $4.163/3.991$ $1.352/5.674$ $-0.152/1.066$ $0.9$ $0.833$ $0.667$
$\text{LLF}_{1}$ (Cross-Over bitrates) $-20.43/15.853$ $4.361/3.702$ $1.342/5.154$ $-0.124/0.84$ $\mathbf{0.967}$ $\mathbf{0.933}$ $\mathbf{0.9}$
@@
TABLE X: Mean and standard deviations of BD-metrics, and closeness of each model’s predicted per-shot quality ladders against fixed and reference bitrate ladders on the validation and test datasets.

### v2

@@ IV-F Predicting Quality or Bitrate using VIF Features
IV-G Predicting Quality or Bitrate using an Ensemble of Low-Level features and VIF features
@@
Fig. 11 shows the performances of the resulting learned bitrate and quality prediction models. It may be observed that the quality prediction models were slightly more successful (better correlations) than the bitrate prediction models. The regressors trained on $\text{LLF}_{2}$$\text{LLF}_{3}$ and $\text{VIFF}_{3}/\text{VIFF}_{6}/\text{VIFF}_{9}$ yield slightly better performance than the other models across resolutions. Similar to the performances of regressors trained on VIF features, regressors trained on low-level features and spatio-temporal VIF features delivered better performance than regressors trained on low-level features and only spatial VIF features. The performances of the best performing models was close to those models trained on either low-level features or VIF features. However, it should be understood that these correlations do not indicate the potential performance gains of these models when used to construct bitrate or quality ladders, or their BD-Rate or BD-VMAF gains, against fixed bitrate ladders.
@@
IV-H Correction Algorithms
In the preceding, we have studied various techniques for predicting per-shot bitrate and quality ladders, using video quality and bitrate prediction models, respectively. Given that machine learning models are prone to errors, it is important to consider corrective measures on the predicted ladders to rectify any inaccuracies. In [Bitrate-Ladder-Construction-using-Visual-Information-Fidelity], we suggested a correction mechanism such that traversing the bitrate ladder from top to bottom (from higher to lower bitrates), will satisfy the condition that the optimal resolution at a current bitrate step should no larger than the optimal resolution at the previous step. This strategy works because:
@@ IV-H Correction Algorithms
  • When constructing a bitrate ladder, we use the predicted quality scores of compressed videos at each bitrate, but over all resolutions, where the resolution having the highest predicted quality is selected as the optimal resolution for encoding at that bitrate.
@@
Tables VII(b) and VIII(b) depict simulated examples of these Top-to-Bottom and Bottom-to-Top ladder correction strategies applied to bitrate ladders and quality ladders, respectively. These correction techniques are applied after the bitrate or quality ladders are constructed.
TABLE VII: An example of Top-to-Bottom bitrate ladder correction.
(a) Before Correction
@@
TABLE VIII: An example of Bottom-to-Top quality ladder correction.
(a) Before Correction
@@
IV-I Bjontegaard Delta Metrics
One of the most commonly used techniques for evaluating or comparing the compression efficiency of video codecs, presets, or encoding modes is by calculating bitrate and quality savings using the Bjøntegaard Delta (BD) method [BD]. It measures the average difference between two RQ curves by fitting a third-order cubic polynomial through the data points, and then calculating the integral to estimate the average bitrate savings, referred to as BD-Rate, and the average quality savings, referred to as BD-Quality (e.g., BD-PSNR or BD-VMAF). Although there are many implementations of this concept, based on the results in [BD-Study], we used the implementation proposed in [BD-Metric].
Since we are performing constant-quality encoding, at each step $b_{i}$ in the bitrate ladder and its corresponding optimal resolution $R_{i}$, to construct the convex hull, we utilize the RQ points of the video from the dataset having resolution of $R_{i}$, with bitrates lying within the range $b_{i}$, $b_{i+1})$. We use this process, rather than selecting the nearest point, to obtain a precise convex hull for BD metric calculations. We applied a similar procedure when constructing convex hulls for quality ladders. We calculated the BD-Rate and BD-VMAF on each video in the validation and test datasets against the fixed bitrate ladder [Fixed-Bitrate-Ladder] and the reference bitrate ladder. The reference bitrate ladders were constructed by sampling the convex hull constructed by exhaustive encoding using the bitrate ladder steps we previously mentioned. We report the mean and standard deviation of BD-Rate and BD-VMAF for each of the compared techniques, on the combined validation and test datasets.
@@ IV-I Bjontegaard Delta Metrics
IV-J Constructing Per-Shot Bitrate and Quality Ladders
TABLE IX: Means and standard deviations of BD-metrics, and closeness of each model’s predicted per-shot bitrate ladders against fixed and reference bitrate ladders on the validation and test datasets. Bitrate ladders have the following bitrates (in kbps) as steps: [500, 1000, 2000, 3000, 4000, 5000, 6000, 7000, 8000, 9000, 10500, 12000, 15000].
@@ IV-J Constructing Per-Shot Bitrate and Quality Ladders
$\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 9.256/20.964-9.256/20.964 $2.084/4.369$ $14.053/11.451$ 2.144/1.933-2.144/1.933 $0.500$ $0.433$ $0.167$
$\text{LLF}_{2}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.429/16.747-20.429/16.747 $4.323/4.092$ $0.439/4.791$ 0.016/1.004-0.016/1.004 $0.967$ $0.933$ $0.767$
$\text{VIFF}_{1}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 12.238/22.332-12.238/22.332 $2.790/4.704$ $9.780/13.017$ 1.519/2.145-1.519/2.145 $0.700$ $0.600$ $0.400$
$\text{VIFF}_{2}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 16.022/20.536-16.022/20.536 $3.521/4.678$ $5.381/8.955$ 0.922/1.617-0.922/1.617 $0.767$ $0.767$ $0.633$
$\text{VIFF}_{3}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.372/18.170-17.372/18.170 $3.665/4.184$ $3.746/5.768$ 0.628/1.039-0.628/1.039 $0.767$ $0.700$ $0.567$
$\text{VIFF}_{4}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 16.852/18.941-16.852/18.941 $3.624/4.184$ $4.613/7.811$ 0.717/1.253-0.717/1.253 $0.767$ $0.700$ $0.567$
$\text{VIFF}_{5}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.822/18.324-17.822/18.324 $3.853/4.243$ $3.396/7.565$ 0.463/1.213-0.463/1.213 $0.800$ $0.767$ $0.633$
$\text{VIFF}_{6}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.655/18.778-17.655/18.778 $3.756/4.275$ $3.661/6.338$ 0.628/1.089-0.628/1.089 $0.767$ $0.700$ $0.600$
$\text{VIFF}_{7}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 16.596/17.158-16.596/17.158 $3.618/4.013$ $5.186/8.562$ 0.659/1.278-0.659/1.278 $0.833$ $0.767$ $0.567$
$\text{VIFF}_{8}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.308/17.842-17.308/17.842 $3.714/4.174$ $4.250/7.433$ 0.637/1.289-0.637/1.289 $0.800$ $0.700$ $0.567$
$\text{VIFF}_{9}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.764/17.275-17.764/17.275 $3.689/3.916$ $3.529/6.726$ 0.525/1.164-0.525/1.164 $0.800$ $0.767$ $0.600$
$\text{LLF}_{2}$, $\text{VIFF}_{1}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.151/16.703-20.151/16.703 $4.195/3.889$ $1.025/4.703$ 0.086/0.933-0.086/0.933 $0.967$ $0.867$ $0.733$
$\text{LLF}_{2}$, $\text{VIFF}_{2}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.523/17.207-20.523/17.207 $4.390/4.071$ $0.829/5.388$ 0.065/1.073-0.065/1.073 $0.967$ $0.867$ $0.767$
$\text{LLF}_{2}$, $\text{VIFF}_{3}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.201/16.083-20.201/16.083 $4.259/3.850$ $1.113/5.015$ 0.089/0.943-0.089/0.943 $1.000$ $0.900$ $0.800$
$\text{LLF}_{2}$, $\text{VIFF}_{4}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.314/16.231-20.314/16.231 $4.327/3.981$ $1.005/5.852$ 0.051/0.987-0.051/0.987 $0.967$ $0.933$ $0.833$
$\text{LLF}_{2}$, $\text{VIFF}_{5}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.379/16.601-20.379/16.601 $4.283/3.864$ $0.764/4.318$ 0.047/0.846-0.047/0.846 $0.933$ $0.867$ $0.800$
$\text{LLF}_{2}$, $\text{VIFF}_{6}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ $\textbf{20.688}/16.624$ $\textbf{4.425}/4.032$ $\textbf{0.347}/4.544$ $\textbf{0.010}/0.916$ $0.967$ $0.933$ $0.833$
$\text{LLF}_{2}$, $\text{VIFF}_{7}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.473/16.738-20.473/16.738 $4.288/3.943$ $0.824/4.721$ 0.080/0.979-0.080/0.979 $0.967$ $0.867$ $0.667$
$\text{LLF}_{2}$, $\text{VIFF}_{8}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.373/17.039-20.373/17.039 $4.374/4.016$ $0.794/3.809$ 0.067/0.756-0.067/0.756 $0.967$ $0.833$ $0.767$
$\text{LLF}_{2}$, $\text{VIFF}_{9}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.086/16.797-20.086/16.797 $4.267/4.095$ $1.037/5.256$ 0.092/1.008-0.092/1.008 $0.967$ $0.800$ $0.733$
$\text{LLF}_{1}$ (Cross-Over bitrates) 20.430/15.853-20.430/15.853 $4.361/3.702$ $1.342/5.154$ 0.124/0.840-0.124/0.840 $\mathbf{0.967}$ $\mathbf{0.933}$ $\mathbf{0.900}$
The construction of per-shot reference bitrate ladders is achieved by sampling the convex hulls computed on videos from both the validation and test datasets. The reference bitrate ladder demonstrated a BD-Rate performance having a mean improvement of $20.63\%$ (negative BD-rate means a gain in bitrate) and a standard deviation of $17.124\%$, and a BD-VMAF performance having a mean improvement of $4.473$ (positive means gain in quality) and standard deviation $4.139$ against the fixed bitrate ladder [Fixed-Bitrate-Ladder]. Our models were compared against techniques that predict cross-over bitrates or cross-over VMAF points. For more accurate comparisons, we computed the performance of bitrate and quality ladders constructed using true cross-over bitrates and VMAFs points, respectively. We found that the bitrate ladder constructed using true cross-over bitrates, attained a BD-rate performance having a mean improvement of $21.653\%$, and a standard deviation of $16.835\%$. It also obtained a BD-VMAF performance having a mean improvement of $4.712$ and a standard deviation of $4.091$. In a similar vein, the quality ladder constructed using true cross-over VMAF points, demonstrated a BD-rate performance having a mean improvement of $21.847\%$ and a standard deviation of $16.254\%$, and a BD-VMAF performance having a mean improvement of $4.752$ and standard deviation $4.145$.
@@
TABLE X: Means and standard deviations of BD-metrics, and closeness of each model’s predicted per-shot quality ladders against fixed and reference bitrate ladders on the validation and test datasets. Quality ladders have the following VMAF scores as steps: [25, 35, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 92.5].